Vo sin 2π ft rmsA 93.0-mH inductor and a 6.50-μF capacitor are connected in series with a generator whose frequency is 549 Hz. The rms voltage across the capacitor is 2.50 V. Determine the rms voltage across the inductor.
Answer #1

xL = 2*3.14*f*L
= 2*3.14*549*93*10^-3
= 320.6 ohm

xC = 1/(2*3.14*f*C)
= 1/(2*3.14*549*6.50*10^-6)
= 1/(22.4*10^-3)
= 44.6 ohm

given,
vC = Irms*xC
2.5 = Irms*44.6
Irms = 0.0561 Amp

vL = Irms*xL
= 0.0561*320.6
= 18.0 V

Answer: 18.0 V